# Word length predicts word order: “Min-max”-ing drives language evolution

### v2

@@ Abstract
A fundamental concern in linguistics has been to understand how languages change, such as in relation to word order. Since the order of words in a sentence (i.e. the relative placement of Subject, Object, and Verb) is readily identifiable in most languages, this has been a productive field of study for decades (see Greenberg 1963; Dryer 2007; Hawkins 2014). However, a language’s word order can change over time, with competing explanations for such changes (Carnie & Guilfoyle 2000; Crisma & Longobardi 2009; Martins & Cardoso 2018; Dunn et al. 2011; Jäger & Wahle 2021). This paper proposes a general universal explanation for word order change based on a theory of communicative interaction (the Min-Max theory of language behavior) in which agents seek to minimize effort while maximizing information. Such an account unifies opposing findings from language processing (Piantadosi et al. 2011; Wasow 2022; Levy 2008) that make different predictions about how word order should be realized crosslinguistically. The marriage of both “efficiency” and “surprisal” approaches under the Min-Max theory is justified with evidence from a massive dataset of 1,942 language corpora tagged for parts of speech (Ring 2025), in which average lengths of particular word classes correlates with word order, allowing for prediction of basic word order from diverse corpora. The general universal pressure of word class length in corpora is shown to give a stronger explanation for word order realization than either genealogical or areal factors, highlighting the importance of language corpora for investigating such questions.
@@ 1 Introduction
There are two major approaches to studying word order. The first (typological) approach is to observe word orders in natural languages and identify patterns whereby the orders of certain constituents are correlated with other grammatical patterns. This approach has developed multiple crosslinguistic databases with expert determinations of word order (WALS, Dryer & Haspelmath 2013; AUTOTYP, Bickel et al. 2023; Grambank, Skirgård et al. 2023) and has identified statistical universals in which the orders of certain elements increase the likelihood for other patterns to be present in a language (Greenberg 1963; Dryer 2007, 1992, 2011; Verkerk et al. 2025). The second (processing) approach focuses on experimental evidence for how constituents in particular languages are processed in the brain. This approach has identified potential universal tendencies or preferences in processing, notably a preference for “agents first” (Culbertson et al. 2012; Futrell et al. 2015a; Sauppe et al. 2023), “efficiency” (Zipf, 1949; Wasow, 2002; Futrell et al., 2015b), and “surprisal” (Hale 2001; Levy 2008).
@@
While the typological observations do not make predictions about a language’s basic word order (the universals take basic word orders as a starting point and make further predictions, see Verkerk et al. 2025), the processing research does imply some particular claims and predictions about how words should be ordered in languages of the world. For example, the “agents first” (or “Subject first”) bias found in experiments suggests that there is a universal psychological pressure for speakers of all languages to produce agent arguments (“do-ers” of actions, typically represented by animate nouns and pronouns) before other kinds of constituents in sentences. This aligns with observations of the dominant crosslinguistic word order being Verb-final or Subject-initial (Schouwstra & de Swart 2014; Schouwstra et al. 2022).
@@
At the same time, findings from language processing studies suggest that other pressures mediate the realization of word order patterns, namely in relation to “efficiency” and “surprisal”. The efficiency literature largely follows from Zipf’s (1932; 1949) observations that a) languages show a logarithmic distribution whereby the most frequently used words occur much more often than the next most frequent, and that b) the most frequent words in a language are also the shortest. The observation that shorter (or less complex) elements are more likely to occur initially in a sentence, while longer (or more complex) elements occur later, can be traced to Behaghel (1909), and the general idea is that this supports efficient communication. In contrast to this view, the surprisal literature focuses on how salient or predictable items are, suggesting that longer (or more complex) words are more likely to occur first in a sentence since they improve predictability of following constituents (Hale 2001; Levy 2008; Sauppe et al. 2021). This is supported by clinical evidence (Rezaii et al. 2023) and information theoretic studies (Zaslavsky 2020; Tucker et al. 2025).
@@
The lack of testing highlights the complexity of investigating pressures that influence the evolution and realization of language structures, as well as the dearth of datasets and databases that could address these questions. [^1] Datasets that have been used by previous research include the Universal Dependencies Treebanks (Zeman et al. 2024) and ODIN (Xia et al. 2016) projects. The UDT contains 160+ hand-annotated languages available for crosslinguistic investigation, but these are not parallel texts and the number of sentences in each language is variable, with some having fewer than 200 sentences in the v2.14 release. The ODIN project (and its re-packaged form via LECS: https://huggingface.co/datasets/lecslab/glosslm-corpus) has interlinear glossed text (IGT) from over 1,800 languages, but the majority of these are single-word/line examples, and only 300 of the languages in their dataset have more than 100 sentences. Additional cross-linguistic investigations using tagged corpora include Hahn et al. (2020: 51 languages), Jing et al. (2021: 71 languages), and Hahn et al. (2022: 80 languages). For each of these studies the majority of languages are from the Indo-European language family. Besides limiting generalizability of findings, such data risks bias toward a particular group of languages (Blasi et al. 2022), in much the same way that other scientific findings have historically been biased toward WEIRD demographics (Henrich et al. 2010). In particular, questions regarding the length of constituents in relation to their order in sentences requires a large number of sentences tagged for parts of speech (at minimum). Given that languages develop over time via cultural evolution, it is also difficult to separate general cognitive pressures from historical (genealogical) or areal factors. This means that inferring crosslinguistic or universal principles necessitates both a large number of languages as well as sufficient diversity of languages to establish statistical validity.
@@
The current paper takes advantage of a massive tagged parallel corpus to show how these pressures are realized in actual language data in relation to word order. While the dataset is still under development, the baseline data is tagged for parts of speech, which allows us to derive a proxy measure for word order from individual language corpora, as well as to measure other linguistic properties such as the average length of various word types. Results show that basic word order in intransitive clauses is highly correlated with the length of particular word classes, allowing for the prediction of a language’s word order based on corpora not in the dataset and explaining variance better than both genealogy and areal factors.
In the remainder of this paper I first highlight the properties of the dataset (§2). I then discuss three studies that investigate word length in relation to word order (§2.1), the predictive validity of word lengths for word order (§2.2), and the variance predicted by word length in regression models (§2.3). I conclude with a discussion of the implications of these studies in relation to language evolution (§3), presenting the Min-Max theory of language behavior (§3.1) as well as some limitations and future directions regarding the current effort (§3.2).
@@ 2 Methodology and hypotheses
The taggedPBC (Ring 2025) [^2] https://github.com/lingdoc/taggedPBC - see fn 5 for a link to a minimal OSF repository that allows for replication of the studies in this paper. contains pos-tagged parallel text data in 1,942 languages (roughly 25% of the world’s languages), with over 1,800 parallel verses for more than 1,850 of those languages. The dataset represents 155 language families and 78 isolates, dwarfing previously available resources, and the tagging methodology (via crosslingual tag transfer) has been shown to correspond well with existing automated taggers for high-resource languages (as found in the SpaCy and TranKit NLP libraries). [^3] https://spacy.io ; https://nlp.uoregon.edu/trankit Additionally, a novel measure extracted from individual corpora in this dataset (the “N1 ratio”) is significantly correlated (p < 0.001) with expert determinations of a language’s basic word order for intransitive clauses (“SV”, “VS”, “free”; Figure 2) as identified in the typological databases WALS (Dryer & Haspelmath 2013), Grambank (Skirgård et al. 2023), and Autotyp (Bickel et al. 2023). [^4] The procedure for determing the N1 ratio is essentially to count the number of sentences in a corpus that have both arguments (nouns, pronouns, etc) and predicates (verbs, auxiliaries), and then to derive a ratio from the number of argument-first sentences divided by the number of predicate-first sentences in that set. As an example, English [ISO 639-3: eng] has an N1 ratio of 4.51, whereas Irish [ISO 639-3: gle] has an N1 ratio of 1.24, based on their respective corpus in the taggedPBC. This ratio can then be related to other aspects of language, such as word order classifications. See Ring (2025) for further details. What this means is that using this dataset we can derive word order and word length information in a systematic manner, allowing for an investigation of word length in relation to word order.
Figure 2: The N1 ratio and word order in 3 typological databases (p < 0.001; Ring 2025)
@@
The following studies focus on intransitive word order (i.e. the basic word order of intransitive clauses, in which a predicate requires a single “S” argument) in languages found in the taggedPBC. [^5] Minimal data and Python code to replicate the results in this paper can be found at https://osf.io/nsaq2 (3MB). More comprehensive data and code (7GB) can be found at the taggedPBC github repository (fn 2). The first study measures word classes, showing that average lengths of nouns and verbs in a language’s corpus are highly correlated with SV, VS, and free word order classifications when frequency is considered. The second study investigates whether this effect occurs outside of the taggedPBC, showing that information on average lengths of nouns and verbs computed from different corpora also allow for the classification of languages as SV or VS, differentiating between historical and modern languages for which word order changes are known. The third study examines whether the lengths of nouns and verbs in corpora provides a better indication of word order than known areal or genealogical signals, and shows that in heirarchical linear regressions, word length is indeed a better predictor of word order.
@@ 2.1 Study 1: Word length and word order
Study 1 follows previous studies in using word length as a measure of word complexity (Lewis & Frank 2016; Bentz et al. 2016, 2017; Baerman et al. 2015, 2017; Bentz et al. 2023; and see Piantadosi et al., 2011; Meylan & Griffiths, 2021; Koplenig et al., 2022 for further discussion in relation to frequency in corpora), comparing average lengths of particular classes of words (nouns/arguments and verbs/predicates) with word order in individual languages. The two processing claims in the literature (from “efficiency” and “surprisal” approaches) give rise to the following two competing hypotheses regarding this relationship:
@@ 2.1.1 Word order
Of the languages in the taggedPBC, 1,152 are also found in typological databases with categorizations for word order. For the remainder, the N1 ratio was used to impute the (SV, VS, free) category of the remaining 790 languages in the taggedPBC not categorized by WALS, AUTOTYP or Grambank, and the final categorizations were stored for all 1,942 languages. [^6] I trained a Gaussian Naive Bayes classifier on the known data to predict the category (SV, VS, free) of languages for which word order status was unknown, based on the N1 ratio as derived from the taggedPBC. Code to reproduce these results can be found at the linked repository (fn 2). This resulted in 1,548 languages classified as “SV”, 331 languages classified as “VS”, and 63 languages classified as “free”.
2.1.2 Nouns and verbs
To measure word length, I counted the number of characters in a word tagged as Noun/Argument or Verb/Predicate [^7] Here I conducted several tests to see if there were different effects for different word types and groupings. “Argument” is a grouping of words tagged as NOUN, PRON, PROPN in the universal tag set, while “predicate” refers to a grouping of VERB and AUX tags. and computed an average length for each class on a per-language basis. [^8] For example, English [ISO 639-3: eng] arguments have an average length of 5.9 characters, while English predicates have an average length of 5.75. English nouns have an average length of 5.9 characters, while English verbs have an average length of 5.8 characters. Irish [ISO 639-3: gle] arguments have an average length of 6.31 characters, while Irish predicates have an average length of 6.29 characters. Irish nouns have an average length of 5.2 characters, while Irish verbs have an average length of 7.9 characters. In the taggedPBC the average number of (unique) arguments per language is 741, and the average number of predicates is 498. The mean length of arguments is 6.71 characters, and the mean length of predicates is 6.82 characters. Conducting a Student’s T-Test showed that this is statistically significant (p < 0.001), indicating that crosslinguistically, arguments are shorter than predicates. All statistics showed a normal distribution.
@@ 2.1.3 Relationship between word order and noun/verb lengths
To see whether word order is related to the length of nouns and verbs I conducted a repeated-measures ANOVA with the average lengths of nouns and verbs as the repeated measures and the word order (SV, VS, free) as the between subject factors. Analysis of the results of word order and its relation to length of the different argument and predicate categories is quite striking. First, both SV and VS languages have a highly significant effect of word length, [^9] Additional statistical comparisons and plots are available at the linked repository (Fn 2). whereby arguments are, overall, shorter than predicates crosslinguistically (Figure 3; p < 0.001). [^10] It is also worth noting here that in this corpus words are overall shorter in SV languages than VS languages, which in turn are shorter than those in ‘free’ languages. This displays a general observed trend that languages with free word order tend to have more morphology to identify argument roles and other grammatical relations (Koplenig et al. 2017; Yadav et al. 2020).
Figure 3: Normalized lens of Args, Preds, by word order
Figure 4: Frequency-weighted lens of Ns, Vs, by word order
If we consider only nouns and only verbs, the findings are similar when we normalize the dataset by reducing the vocabulary to unique items. However, if we allow for frequency effects (i.e. we do not reduce the vocabulary to unique items) there is a clear correlation with word order: nouns are longer than verbs in SV languages, but shorter than verbs in VS languages (Figure 4; p < 0.001). This is a major finding that we will return to in the discussion. Setting it aside for now, we turn to two different tests of this measure relating to diachrony: predicting word order change, and the potential interaction with family membership (genealogy/descent) and area (geography).

### v3

@@ Abstract
A fundamental concern in linguistics has been to understand how languages change, such as in relation to word order. Since the order of words in a sentence (i.e. the relative placement of Subject, Object, and Verb) is readily identifiable in most languages, this has been a productive field of study for decades (see Greenberg 1963; Dryer 2007; Hawkins 2014). However, a language’s word order can change over time, with competing explanations for such changes (Carnie & Guilfoyle 2000; Crisma & Longobardi 2009; Martins & Cardoso 2018; Dunn et al. 2011; Jäger & Wahle 2021). This paper proposes a general universal explanation for word order change based on a theory of communicative interaction (the Min-Max theory of language behavior) in which agents seek to minimize effort while maximizing information. Such an account unifies opposing findings from language processing (Piantadosi et al. 2011; Wasow 2022; Levy 2008) that make different predictions about how word order should be realized crosslinguistically. The marriage of both “efficiency” and “surprisal” approaches under the Min-Max theory is justified with evidence from a massive dataset of 1,942 language corpora tagged for parts of speech (Ring 2026), in which average lengths of particular word classes correlates with word order, allowing for prediction of basic intransitive word order from diverse corpora. The general universal pressure of word class length in corpora is shown to give a stronger explanation for word order realization than either genealogical or areal factors, highlighting the importance of language corpora for investigating such questions
@@ 1 Introduction
There are two major approaches to studying word order. The first (typological) approach is to observe word orders in natural languages and identify patterns whereby the orders of certain constituents are correlated with other grammatical patterns. This approach has developed multiple crosslinguistic databases with expert determinations of word order (such as WALS, Dryer & Haspelmath 2013; AUTOTYP, Bickel et al. 2023; Grambank, Skirgård et al. 2023) and has identified statistical universals in which the orders of certain elements increase the likelihood for other patterns to be present in a language (Greenberg 1963; Dryer 2007, 1992, 2011; Verkerk et al. 2025). The second (processing) approach focuses on experimental evidence for how constituents in particular languages are processed in the brain. This approach has identified potential universal tendencies or preferences in processing, notably a preference for “agents first” (Culbertson et al. 2012; Futrell et al. 2015a; Sauppe et al. 2023), “efficiency” (Zipf, 1949; Wasow, 2002; Futrell et al., 2015b), and “surprisal” (Hale 2001; Levy 2008).
@@
While the typological observations do not make predictions about a language’s basic word order (the statistical universals take basic word orders as a starting point and make further predictions, see Verkerk et al. 2025), the processing research does imply some particular claims and predictions about how words should be ordered in languages of the world. For example, the “agents first” (or “Subject first”) bias found in experiments suggests that there is a universal psychological pressure for speakers of all languages to produce agent arguments (“do-ers” of actions, typically represented by animate nouns and pronouns) before other kinds of constituents in sentences. This aligns with observations of the dominant crosslinguistic word order being Verb-final or Subject-initial (Schouwstra & de Swart 2014; Schouwstra et al. 2022).
@@
At the same time, findings from additional language processing research suggests that other pressures mediate the realization of word order patterns, namely in relation to “efficiency” and “surprisal”. The efficiency literature largely follows from Zipf’s (1932; 1949) observations that a) languages show a logarithmic distribution whereby the most frequently used words occur much more often than the next most frequent, and that b) the most frequent words in a language are also the shortest. The finding that shorter (or less complex) elements are more likely to occur initially in a sentence, while longer (or more complex) elements occur later, is one that can be traced to Behaghel (1909), and the general idea is that this supports efficient communication. In contrast to this view, the surprisal literature focuses on how salient or predictable items are, suggesting that longer (or more complex) words are more likely to occur first in a sentence since they improve predictability of following constituents (Hale 2001; Levy 2008; Sauppe et al. 2021). This is supported by clinical evidence (Rezaii et al. 2023) and information theoretic studies (Zaslavsky 2020; Tucker et al. 2025).
@@
The lack of testing highlights the complexity of investigating pressures that influence the evolution and realization of language structures, as well as the dearth of datasets and databases that could address these questions. [^1] Datasets that have been used by previous research include the Universal Dependencies Treebanks (UDT; Zeman et al. 2024) and ODIN (Xia et al. 2016) projects. The UDT contains 160+ hand-annotated languages available for crosslinguistic investigation, but these are not parallel texts and the number of sentences in each language is variable, with some having fewer than 200 sentences in the v2.14 release. The ODIN project (and its re-packaged form via LECS: https://huggingface.co/datasets/lecslab/glosslm-corpus) has interlinear glossed text (IGT) from over 1,800 languages, but the majority of these are single-word/line examples, and only 300 of the languages in their dataset have more than 100 sentences. Additional cross-linguistic investigations using tagged corpora include Hahn et al. (2020: 51 languages), Jing et al. (2021: 71 languages), and Hahn et al. (2022: 80 languages). For each of these studies the majority of languages are from the Indo-European language family. Besides limiting generalizability of findings, such data risks bias toward a particular group of languages (Blasi et al. 2022), in much the same way that other scientific findings have historically been biased toward WEIRD demographics (Henrich et al. 2010). In particular, questions regarding the length of constituents in relation to their order in sentences requires a large number of sentences tagged for parts of speech (at minimum). Given that languages develop over time via cultural evolution, it is also difficult to separate general cognitive pressures from historical (genealogical or areal) factors. This means that inferring crosslinguistic or universal principles necessitates both a large number of languages as well as sufficient diversity of languages to establish statistical validity.
@@
The current paper takes advantage of a massive tagged parallel corpus to show how these pressures are realized in actual language data in relation to word order. While the dataset is still under development, the baseline data is tagged for parts of speech, which allows us to derive a proxy measure for word order from individual language corpora. It also allows for measurement of other linguistic properties such as the average length of various word types, while also controlling for potential genealogical and areal bias. Results show that basic word order in intransitive clauses is highly correlated with the length of particular word classes, allowing for the prediction of a language’s word order based on corpora not in the dataset and explaining variance better than both genealogy and areal factors.
In the remainder of this paper I first highlight the properties of the dataset (§2). I then present three studies that investigate word length in relation to word order (§2.1), the predictive validity of word lengths for word order (§2.2), and the variance predicted by word length in regression models over and above language family and linguistic area (§2.3). I conclude with the implications of these studies in relation to language evolution (§3), discussing how the Min-Max theory of language behavior unites competing pressures (§3.1), as well as some limitations and future directions regarding the current effort (§3.2).
@@ 2 Methodology and hypotheses
The taggedPBC (Ring 2026) [^2] https://github.com/lingdoc/taggedPBC - see fn 5 for a link to a minimal OSF repository that allows for replication of the studies in this paper. contains pos-tagged parallel text data in 1,942 languages (roughly 25% of the world’s languages), with over 1,800 parallel verses for more than 1,850 of those languages. The dataset represents 155 language families and 78 isolates, dwarfing previously available resources, and the tagging methodology (via crosslingual tag transfer) has been shown to correspond well with existing automated taggers for high-resource languages (as found in the SpaCy and TranKit NLP libraries). [^3] https://spacy.io ; https://nlp.uoregon.edu/trankit Additionally, a novel measure extracted from individual corpora in this dataset (the “N1 ratio”) significantly correlates (p < 0.001) with expert determinations of a language’s basic word order for intransitive clauses (“SV”, “VS”, “free”; Figure 2) as identified in the typological databases WALS, Grambank, and AUTOTYP. [^4] The procedure for determing the N1 ratio is essentially to count the number of sentences in a corpus that have both arguments (nouns, pronouns, proper nouns) and predicates (verbs, auxiliaries), and then to derive a ratio from the number of argument-first sentences divided by the number of predicate-first sentences in that set. As an example, English [ISO 639-3: eng] has an N1 ratio of 4.51, whereas Irish [ISO 639-3: gle] has an N1 ratio of 1.24, based on their respective corpus in the taggedPBC. This ratio can then be related to other aspects of language, such as word order classifications. See Ring (2026) for further details. What this means is that using this dataset we can derive word order and word length information in a systematic manner, allowing for an investigation of word length in relation to word order.
Figure 2: The N1 ratio and word order in 3 typological databases (p < 0.001; Ring 2026)
@@
The following studies focus on intransitive word order (i.e. the basic word order of intransitive clauses, in which a predicate requires a single “S” argument) in languages found in the taggedPBC. [^5] Minimal data and Python code to replicate the results in this paper can be found at https://osf.io/nsaq2 (3MB). More comprehensive data and code (7GB) can be found at the taggedPBC github repository (fn 2). The first study (§2.1) measures word classes, showing that average lengths of nouns and verbs in a language’s corpus are highly correlated with SV, VS, and free word order classifications when frequency is considered, and that this holds even when controlling for family and areal bias. The second study (§2.2) investigates whether this effect occurs outside of the taggedPBC, showing that information on average lengths of nouns and verbs computed from different corpora also allow for the classification of languages as SV or VS, differentiating between historical and modern languages for which word order changes are known. The third study (§2.3) examines whether the lengths of nouns and verbs in corpora provides a better indication of word order than known areal or genealogical signals, and shows that in heirarchical linear regressions, word length is indeed a better predictor of word order.
@@ 2.1 Study 1: Word length and word order
Study 1 follows previous research in using word length as a measure of word complexity (Lewis & Frank 2016; Bentz et al. 2016, 2017; Baerman et al. 2015, 2017; Bentz et al. 2023; and see Piantadosi et al., 2011; Meylan & Griffiths, 2021; Koplenig et al., 2022 for further discussion in relation to frequency in corpora), comparing average lengths of particular classes of words (nouns/arguments and verbs/predicates) with word order in individual languages. [^6] In the following studies romanized orthographic characters are used as a proxy for phonemes. The two processing claims in the literature (from “efficiency” and “surprisal” approaches) give rise to the following two competing hypotheses regarding this relationship:
@@ 2.1.1 Word order
Of the languages in the taggedPBC, 1,152 are also found in typological databases with categorizations for word order. For the remainder, the N1 ratio was used to impute the (SV, VS, free) category of the remaining 790 languages in the taggedPBC not categorized by WALS, Grambank or AUTOTYP, and the final categorizations were stored for all 1,942 languages. [^7] I trained a Gaussian Naive Bayes classifier on the known data to predict the category (SV, VS, free) of languages for which word order status was unknown, based on the N1 ratio as derived from the taggedPBC. Code to reproduce these results can be found at the linked repository (fn 2). This resulted in 1,548 languages classified as “SV”, 331 languages classified as “VS”, and 63 languages classified as “free”.
2.1.2 Measuring word length by class
To measure word length, I counted the number of characters in a word tagged as Noun/Argument or Verb/Predicate [^8] Here I conducted several tests to see if there were different effects for different word types and groupings, where “argument” is a grouping of words tagged as NOUN, PRON, PROPN in the universal tag set, while “predicate” refers to a grouping of VERB and AUX tags. and computed an average length for each class on a per-language basis. [^9] For example, English [ISO 639-3: eng] arguments have an average length of 5.9 characters, while English predicates have an average length of 5.75. English nouns have an average length of 5.9 characters, while English verbs have an average length of 5.8 characters. Irish [ISO 639-3: gle] arguments have an average length of 6.31 characters, while Irish predicates have an average length of 6.29 characters. Irish nouns have an average length of 5.2 characters, while Irish verbs have an average length of 7.9 characters. In the taggedPBC the average number of (unique) arguments per language is 741, and the average number of predicates is 498. The mean length of arguments is 6.71 characters, and the mean length of predicates is 6.82 characters. Conducting a Student’s T-Test showed that this is statistically significant (p < 0.001), indicating that crosslinguistically, arguments are shorter than predicates. All statistics showed a normal distribution.
Since we are working with corpus data, there are two different length measures that we can compute for word classes in each language. We can compute average lengths for classes based on 1) the unique items in that class, and 2) the total items in the class. The first measure is similar to the information that can be gleaned from a wordlist or dictionary, where infrequent terms have the same weight as frequent items. The second measure gives more weight to frequent items. Since we know that frequency of occurrence in corpora is a factor in constituent predictability (Piantadosi et al., 2011; Meylan & Griffiths, 2021; Koplenig et al., 2022), it is worth observing whether the two measures differ in terms of their relationship with word order.
@@ 2.1.3 Relationship between word order and noun/verb lengths
To see whether word order is related to the length of nouns and verbs we can first visualize the relationships by plotting the means of word class lengths (for arguments, predicates, nouns, verbs, etc) in relation to word order (SV, VS, free). Figure 3 shows the means of unique nouns and verbs (measure #1 above) in relation to word order, while Figure 4 shows the means of frequency-weighted nouns and verbs in relation to word order (measure #2 above). [^10] Additional statistical comparisons and plots are available at the linked repository (Fn 2). Observationally, the plots indicate that when frequency is taken into account, nouns are longer than verbs in SV languages but shorter than verbs in VS languages. [^11] It is also worth noting here that in this corpus words are overall shorter in SV languages than VS languages, which in turn are shorter than those in ‘free’ languages. This displays a general observed trend that languages with free word order tend to have more morphology to identify argument roles and other grammatical relations (Koplenig et al. 2017; Yadav et al. 2020).
Figure 3: Normalized lengths of Nouns, Verbs, by word order
Figure 4: Frequency-weighted lengths of Ns, Vs, by word order
To verify this I conducted a repeated-measures ANOVA with the average lengths of nouns and verbs as the repeated measures and the word order (SV, VS, free) as the between subject factors. Results of the ANOVA indicate that the effect of frequency-weighted noun and verb lengths is highly significant (p < 0.001) for differentiating between word orders. However, the ANOVA does not account for potential effects of language relatedness or geographical area, a subject we turn to next.
2.1.4 Controlling for genealogical and areal bias
Given that human languages develop over time through interaction, it has long been recognized that the evolution of particular linguistic features can potentially be accounted for by descent from a common ancestral state or by contact between languages spoken in the same area (i.e. Galton’s problem, c.f. Bromham 2024). It is therefore common practice to account or control for such genealogical or areal effects in typological studies. While there are several proposals on the best way to handle this (ibid, see also extensive discussion in Guzmán Naranjo & Becker 2022), a basic method is to group observed languages according to their family membership and/or macroarea.
In keeping with such approaches, for each language I identified family membership and linguistic area on the basis of Glottolog classifications, [^12] Glottolog (Hammarström et al. 2024; https://www.glottolog.org) contains expert annotations for language family affiliations for over 7,000 languages, along with ‘macroarea’ identifiers and additional aggregated data from a variety of sources. For family membership I take the highest level node, based on the observation that all modern languages are the same age, and while their pathways/histories may differ, this does not impact whether they can be considered part of the same phylogenetic tree. filling in missing data via Ethnologue and other sources, and included them as grouping criteria in a linear mixed-effects regression model. From the dataset of 1,942 languages I examined four word-length-related continuous features to identify potential effects of genealogy and area on word order. The first two features are the frequency-weighted average length of nouns and verbs in a corpus. The third feature is the difference in length between nouns and verbs, which is related to minimization of complexity, based on the idea that a larger difference between lengths of word classes may effect the likelihood of that particular class being placed first. The fourth feature is the ratio of noun and verb lengths, and is related to maximization of complexity (via “entropy”). Here a lower ratio indicates that the particular word classes are less likely to be easily differentiable (and therefore predictable) based on their length.
For each feature I conducted separate linear regressions, treating both grouping factors (family membership, macroarea) as random effects with word order (SV, VS, free) as the dependent variable and the feature as the independent variable (Table 1). [^13] Tables in this section use the Mixed Linear Model from the Python statsmodels library, and all features are scaled with the MinMaxScaler from scikit-learn (which scales values to between 0 and 1). Other approaches can easily be explored using code at the taggedPBC Github repository. Each feature was found to be significant, with little effect from either family membership or macroarea. Treating family and macroarea as nested within each other did not change this result.
Number of Languages: 1,942
iv Coef. Std.Err. z P>|z|
NVlensdiff 0.281 0.094 2.996 0.003
NVlensratio 0.255 0.080 3.174 0.002
Nlen_freq 0.433 0.109 3.957 0.000
Vlen_freq 0.499 0.100 4.998 0.000
Table 1: Word length features in relation to word order (SV, VS, free) with family/area factors (for 1,942 languages)
As an additional check (Table 2), I filtered the dataset based on number of languages in each family, first excluding single languages and isolates (leaving 1,864 languages), then excluding families with fewer than 10 members (leaving 1,653 languages). In each case all features remained significant, with little effect from family membership or macroarea.
Number of Languages: 1,864
iv Coef. Std.Err. z P>|z|
NVlensdiff 0.321 0.095 3.360 0.001
NVlensratio 0.283 0.081 3.478 0.001
Nlen_freq 0.448 0.111 4.045 0.000
Vlen_freq 0.539 0.101 5.318 0.000
Number of Languages: 1,653
iv Coef. Std.Err. z P>|z|
NVlensdiff 0.382 0.101 3.766 0.000
NVlensratio 0.323 0.086 3.751 0.000
Nlen_freq 0.332 0.122 2.710 0.007
Vlen_freq 0.475 0.108 4.394 0.000
Table 2: Word length features in relation to word order (SV, VS, free) with family/area factors (for 1,864 and 1,653 languages respectively)
It is worth noting here that coefficients of the independent variables are in most cases not above 0.5. However, when we see that these coefficients are combined with z-values above 2.0 and highly significant P-values (all below 0.01), it indicates a robust effect (Luke 2017) for each feature. It is also important to observe that two of the features are derived from the others, which means there is likely to be high collinearity, such that combining their values in a regression results in the most robust feature (’Vlen_freq’ in this case) being found to be significant while the others are not. That features are found to be collinear does not mean that they should be excluded from a model, however - features that point in the same direction can still add extra information, which can be useful in other tasks such as classification, to which we turn next. These observations from linear regressions are simply one way of looking at the data, highlighting the importance of word length in relation to word order. By including family and macroarea as independent random effects the results indicate that this relationship is not biased by either factor.